# Run-Time Monitoring of ERTMS/ETCS Control Flow by Process Mining

### v1

@@ 3.2.2 Online Conformance Checking
Similarly to process discovery, conformance checking can also be performed through a variety of algorithms, such as footprint-based, token-based, and alignment-based algorithms [1]. In this paper, we focus on alignment-based algorithms. These algorithms build alignments, which describe the path across the execution model that either matches or is the best approximation of the sequence of the partial trace. Using the approach in [35], the alignments can be translated into tabular data $\mathcal{D}\in\mathbb{R}^{k\times m}$, where $m$ is the cardinality of the set $\mathcal{A}=\mathcal{A}_{M}\cup\mathcal{A}_{\Sigma}$, i.e., the set of the events contained in $M$ ($\mathcal{A}_{M}$) and present in $\Sigma$ ($\mathcal{A}_{\Sigma}$). In particular, any deviation associated with a particular transition is recorded into a cell within the row of the table associated with the partial trace diagnoses. We call this tuple of diagnosis $d_{\sigma}=\{d_{1},\dots,d_{|\sigma|}\}$. We aim to perform online conformance checking by employing a window-based approach [5], i.e., once a certain number of run-time events are collected from the system, these compose the partial trace and a partial alignment is computed.
@@ 3.3 Anomaly Detection and Explanation
In the third phase, the conformance-checking diagnoses of non-conformant traces are analyzed to spot the occurrence of specific anomalies. Given the nature of data, the detection and explanation — or diagnosis — of anomalies has to be conducted via an unsupervised mechanism, which allows to learn rules from an unlabeled training set [8].
@@ 3.3.1 Clustering
The literature offers a wide variety of algorithms for unsupervised anomaly detection, including, but not limited to [17], [38]: angle-based, neighbour(density)-based, statistical, neural networks, or clustering. Particularly, clustering is a type of unsupervised machine learning that aims at discovering clusters (i.e., subsets of data points, or diagnoses that are similar between each other) in training data, allowing to map a novel data point, or diagnosis $d_{\sigma}$, to a specific cluster, or no cluster at all (i.e., an outlier). Clustering techniques are particularly fitting for this study, as they partition the input space into “groups” that are easy to visualize and that are easier to understand than with other methods, which are not always transparent to the user. Moreover, they could even be used for multi-class classification, either natively [29], or by running a labeling procedure that assigns a label to each cluster that is found by the clustering process.
@@ 3.3.2 Anomaly Explanation
In our study, each cluster is mapped with a custom explanation that relies on the number of misalignments recorded in different steps of the execution flow. Specifically, given a cluster $c$ and the l diagnoses $d_{1},\dots,d_{l}\in c$, the anomaly explanation associated with cluster $c$ is vector $s_{c}\in\mathbb{R}^{m}$ such that $s_{c}(i)=\sum_{j=1}^{l}d_{j}(i)$, where $i$ is the $i$-th activity $a_{i}\in\mathcal{A}$. The set of all anomaly explanations provides the statistics $\mathcal{S}\in\mathbb{R}^{o\times m}$, where $o$ denotes the number of clusters. Moreover, we consider additional quantities that are derived by each explanation $s_{c}$, namely $\mathcal{S}_{c,i}=\sum_{j\in\mathcal{A}_{i}}s_{c}(j),i\in comp$. These provide a detailed view of the explanations for each monitored component compcomp, which has a subset $\mathcal{A}_{i}\subseteq\mathcal{A},i\in comp$ of activities associated with it. Thus, we assign the probability $P(i)=\frac{s_{c,i}}{\sum_{i\in comp}s_{c,i}}$, which denotes the probability that an anomaly occurred in component $i$.
@@ 3.3.3 Labeling
Once these quantities are computed, we are able to automatically label each cluster based on the maximum probability, i.e., given cluster $c$, $label(c)=\max_{i}\{P(i)\}$. In other words, if a cluster contains diagnoses that are more likely to have misalignments in the i-th component, the cluster will be labeled as corresponding to an anomaly in such component. Note that this process of deriving explanations and compute the a-posteriori labeling of a clustering result does not require any knowledge of labels (thus it is entirely unsupervised) and solely relies on knowing which steps of the process are related to a specific component, which is something that can be easily derived by any functional specification of a problem, system, or protocol.
@@ 3.3.4 Anomaly Localization
Once a new trace $\sigma_{test}$ is observed, its diagnosis tuple $d_{\sigma test}$ is derived, statistics $s_{c}\in\mathbb{R}^{m}$ are computed and compared to existing clusters, assigning the trace $\sigma_{test}$ to the closest cluster. Straightforwardly, the trace $\sigma_{test}$ will be labeled as an $i\in comp$ anomaly if the closest cluster was previously labeled as containing diagnosis tuples affected by a fault in the i-th component. This completes the detection and explanation step, which uses diagnoses (non-conformances) of a trace with respect to its expected process flow to detect anomalies, diagnose their root cause, and localize their origin.
@@ 4.1 Behavior Characterization
Behavior characterization first involves simulating the RBC/RBC Handover scenario through a high-level description of its process according to the SUBSET-026 document of the ERTMS standard. Our description compiles the various sequences identified in the document under the chosen scenario into a BPMN diagram. Since the diagram models a workflow from start to end, the model can be simulated and various traces can be collected. After these traces are collected, they are further pre-processed to mimic a realistic monitoring process of the computer-based system. Specifically, we aim to mimic the procedural programming paradigm, which is commonly found in the real implementation of ERTMS/ETCS systems. Using the activities of the traces generated through the BPMN diagram, our approach records the invocation of a procedure each time a new activity is executed as follows. Let $X$ be an ERTMS/ETCS component and $\mathcal{P}_{X}$ be the procedures implemented by $X$. One of the procedures is executed with a given probability for each distinct activity carried out by the component. It is expected that normal behavior follows a given control-flow with a high probability, whereas anomalous behavior involves the alteration of such normal control flow with the improper execution of other procedures. This is enforced using the following probability scheme. Let $\mathcal{A}_{HO}$ be the set of activities of the RBC/RBC Handover scenario, $\mathcal{X}=\{ARBC,HRBC,EVC,RTM\}$ the set of ERTMS/ETCS components involved in the use case, $\mathcal{P}_{X}=\{p_{0},\dots,p_{q}\}$ the $q\in\mathbb{N}$ procedures implemented by $X\in\mathcal{X}$, and $\rho\in[0,1)$ a real number. The procedure execution probability scheme $PEP$ assigns a probability value to a procedure $p\in\mathcal{P}_{X}$ corresponding to $a\in\mathcal{A}_{HO}$ such that
$PEP(a,p,\rho)=\begin{cases}\rho\quad\textrm{if $p==p_{0}$,}\\ \frac{1-\rho}{|\mathcal{P}_{X}|-1}\quad\textrm{otherwise.}\end{cases}$
The motivation behind the first case of the equation is the possibility to simulate traces that are very similar to each other for high values of $\rho$. We generate the execution traces of RBC/RBC Handover as follows.
@@
    Set $|\mathcal{P}_{X}|$ for each component $X$ involved in the RBC/RBC Handover use case;
@@
    For each activity $a\in\mathcal{A}_{HO}$, enforce a random order to the procedures that may be executed for that activity;
@@
    For each activity of the traces of $N_{norm}$, sample a procedure through $PEP$ using $\rho$.
In this preliminary setup, we have decided to set $|\mathcal{P}_{X}|=10$ for each component and generate 100 $N_{norm}$ traces. The dataset generation procedure has been published on the GitHub repository [^3] https://github.com/francescovitale/pm_ertms.
@@
We used the integer linear programming-based miner with 75% tolerance to noise to build a workflow Petri net from $N_{norm}$. Fig. 4 shows the Petri net that the algorithm extracted. The Petri net contains two specific places: the source and the sink. The source place, marked with ‘1’, allows triggering the first transition, whereas the sink place marks the end of the workflow. The Petri net has 21 places, 20 transitions and 78 arcs. The Woflan diagnosis [33] reports that there always exists a path leading to the sink and that there are no dead transitions.
@@ 4.2 Behavior Diagnosis
In this part, we aim to extract online actions from run-time traces and check them against the execution model obtained in the previous step. Run-time traces can be generated with the same procedure above. In particular, we generated 400 $N_{anom}$ traces and injected control-flow anomalies. As there are 4 components, we split the traces into four sets of 100 traces $N_{anom,ARBC}$, $N_{anom,HRBC}$, $N_{anom,EVC}$, and $N_{anom,RTM}$. For each trace of a given set $N_{anom,X}$, we scan the trace procedure-by-procedure. If the procedure is executed by $X$, we randomly inject either a wrongly-ordered, skipped or wrong procedure control-flow anomaly. To account for the randomness of the simulation, we replicated the simulation ten times with different random seeds.
After we obtained the four sets of traces, we proceeded with the extraction of online actions by splitting each trace into multiple traces according to a window length. The number of online actions during this phase depends on the window length. We set three different window lengths, 5, 10 and 15, to evaluate the change in the results of the subsequent steps. For example, if a given trace has 20 traces and the window length is 10, it is split into 2 subtraces, where each subtrace contains the online actions. Regardless of the window length, each subtrace is checked against the execution model to collect the conformance checking diagnoses to use in the subsequent steps.
@@ 4.3 Anomaly Detection and Explanation
The conformance checking diagnoses of the four sets of traces are clustered to identify the different anomalies. In particular, to evaluate the quality of different clustering techniques in correctly classifying the anomalies, we split the diagnoses into a training and a test set. We held out 25% of the diagnoses and used the training set to extract the clusters. Hence, there are 300 training diagnoses and 100 test diagnoses. Each cluster obtained from the training set was labeled with the type of anomaly injected. For example, all the clusters formed with the diagnoses of $N_{anom,ARBC}$ are labeled ARBC. The test diagnoses are subsequently assigned to each cluster and labeled accordingly. For completeness of our analysis, we selected multiple clustering algorithms, ensuring to select candidate clustering algorithms that are as diverse as possible between each other. We selected clustering algorithms that rely (e.g., K-Means) or do not rely on centroids as (H)DBSCAN, algorithms that automatically derive the optimal number of clusters, and some that require the user to input the desired number of clusters as a parameter. Overall, we selected 10 algorithms: BIRCH, K-Means, Mini-Batch KMeans, Ward Hierarchical Clustering, Spectral Clustering, HDBSCAN, DBSCAN, OPTICS, MeanShift, AffinityPropagation, whose implementation is available in the open-source Python library scikit-learn. Algorithms that required the number of clusters as input were exercised 5 times, each with the input parameter of 10, 30, and 50 clusters.
The conformance checking diagnoses collect the misalignments of the 40 possible procedures. These procedures are categorized according to the component that executes them: procedures 0 to 9: ARBC component; procedures 10 to 19: EVC component; procedures 20 to 29: HRBC component; procedures 30 to 39: RTM component, thus $comp=\{ARBC,EVC,HRBC,RTM\}$. For each cluster, the statistics $s_{c}$ will be used to compute the quantities $\mathcal{S}_{c,ARBC}$, $\mathcal{S}_{c,EVC}$, $\mathcal{S}_{c,HRBC}$, $\mathcal{S}_{c,RTM}$, each quantifying the average per-activity misalignment. Consequently, and regardless of the specific clustering algorithm, each cluster is labeled according to the process in Section 3.3.3 either as ARBC, EVC, HRBC or RTM.
@@ 5.1 Evaluation Metrics
There are a wide variety of metrics to measure the classification performance of multi-class classifiers [19], the most common being accuracy, which is the percentage of the correct classifications over all inferences, which is typically used in its balanced accuracy formulation. Other commonly used metrics are those that are primarily meant for binary classification (e.g., precision, recall, F-measure, matthews correlation coefficient) but may be adapted to multi-class scenarios by performing an average per class. For this study, balanced accuracy is enough to quantify the performance of clustering algorithms to detect, explain and diagnose specific faults in the RBC/RBC Handover process.
Then, there is a specific subset of metrics that specifically aim at assessing the performance of clustering processes. Ideally, a cluster should contain all and only instances of a specific class; thus, the following has been defined:
  • A clustering result satisfies homogeneity if all of its clusters contain only data points which are members of a single class.
  • A clustering result satisfies completeness if all the data points that are members of a given class are elements of the same cluster.
Similarly to what F-measure does with precision and recall, V-measure [27] computes a harmonic mean of homogeneity and completeness, providing a [0, 1] metric value which accounts for both aspects.
Summarizing, we quantify the performance of clustering algorithms using (balanced) accuracy and v-measure.
Last, we compute the average per-activity misalignment of each data point with the same logic of cluster-related $\mathcal{S}_{c,ARBC}$, $\mathcal{S}_{c,EVC}$, $\mathcal{S}_{c,HRBC}$, $\mathcal{S}_{c,RTM}$, which use the cluster statistics $s_{c}$ as baseline. At the granularity of data points, or individual misalignments of a trace, we apply the same formulas, but using the misalignmants as baseline instead of the cluster statistics. This results in four distinct values to be computed per data point: $\mathcal{S}_{ARBC}$, $\mathcal{S}_{EVC}$, $\mathcal{S}_{HRBC}$, $\mathcal{S}_{RTM}$.
5.2 Results and discussion

### v2

@@ 3.2.2 Online Conformance Checking
Similarly to process discovery, conformance checking can also be performed through a variety of algorithms, such as footprint-based, token-based, and alignment-based algorithms [aalst2022pmhb]. In this paper, we focus on alignment-based algorithms. These algorithms aim to build an alignment between a trace and the normative model by finding the path across it — the model trace — that either matches or is the best approximation of the trace. Using the approach in [vitale2025cfad], the mismatches between the $k$ actual traces and their corresponding model traces can be translated into “conformance checking diagnoses", i.e., tabular data $\mathcal{D}\in\mathbb{N}^{k\times m}$, where $m$ is the cardinality of the set of activities $\mathcal{A}$. Each row of $\mathcal{D}$ represents the conformance checking diagnosis of a trace $\sigma$: $d_{\sigma}=\{d_{1},\dots,d_{m}\}$, where $d_{i}$ is the number of misalignments associated with activity $a_{i}\in\mathcal{A}$. We aim to perform online conformance checking by employing a window-based approach [burattin2020streamingpdcc], i.e., once a certain number of run-time events are collected from the system, these compose the partial trace and a partial alignment is computed. The partial trace is enlarged incrementally as more windows of activities are collected.
@@ 3.3 Anomaly Detection and Explanation
In the third phase, the conformance-checking diagnoses of non-conformant traces are analyzed to spot the occurrence of specific anomalies. As the data is unlabeled, the detection, explanation, labeling, and localization of anomalies are conducted in an unsupervised manner.
@@ 3.3.1 Clustering
The literature offers a wide variety of algorithms for unsupervised anomaly detection, including, but not limited to [goldstein2016comparative], [zoppi2021unsupervised]: angle-based, neighbour(density)-based, statistical, neural networks, or clustering. Particularly, clustering is a type of unsupervised machine learning that aims at discovering clusters (i.e., subsets of data points, or diagnoses that are similar between each other) in training data, allowing to map a novel data point, or diagnosis $d_{\sigma}$, to a specific cluster, or no cluster at all (i.e., an outlier). Clustering techniques are particularly fitting for this study, as they partition the input space into “groups” that are easy to visualize and easier to understand than with other methods, which are not always transparent to the user. In our methodology, each tuple of diagnoses $d_{\sigma}\in\mathcal{D}$ is assigned to a cluster using one of the techniques above.
@@ 3.3.2 Anomaly Explanation
This step aims to explain the different clusters formed with $\mathcal{D}$, analyzing the misalignments of each tuple belonging to the different clusters. Specifically, given $C=\{c_{1},c_{2},\dots,c_{o}\}$ the set of $o$ clusters and $\mathcal{D}_{c}\subseteq\mathcal{D}$ the set of diagnoses belonging to $c\in C$, the anomaly explanation associated with cluster $c$ is vector $s_{c}\in\mathbb{R}^{m}$ such that $s_{c}(i)=\sum_{j=1}^{l}d_{j}(i)$, where $i$ is the $i$-th activity $a_{i}\in\mathcal{A}$. Next, given $COMP=\{comp_{1},comp_{2},\dots,comp_{p}\}$ the set of $p$ ERTMS/ETCS L2 components of the target scenario and $comp\in COMP$ a specific component, we compute $s_{c,comp}=\sum_{j\in\mathcal{A}_{comp}}s_{c}(j)$, which provides the amount of misalignments of the activities $\mathcal{A}_{comp}\subseteq\mathcal{A}$ linked to compcomp. $s_{c,COMP}=\{s_{c,comp_{1}},s_{c,comp_{2}},\dots,s_{c,comp_{p}}\}$ is the anomaly explanation of cluster $c$. Finally, we calculate the probabilities of each component compcomp for a given cluster $c$ by computing $P(c,comp)=\frac{s_{c,comp}}{\sum_{comp_{i}\in COMP}s_{c,comp_{i}}}$. These values denote the probabilities that an anomaly occurred in the specific components, and are used in the subsequent labeling step.
@@ 3.3.3 Labeling
Once the probabilities of each cluster are computed, we are able to label the clusters based on the maximum probability. Specifically, if a cluster contains diagnoses that are more likely to have misalignments in a specific component, the cluster will be labeled as corresponding to an anomaly in such component. Note that this process of deriving explanations and compute the a-posteriori labeling of a clustering result does not require any knowledge of labels (thus it is entirely unsupervised) and solely relies on knowing which steps of the process are related to a specific component, which is something that can be easily derived by any functional specification of a problem, system, or protocol.
@@ 3.3.4 Anomaly Localization
When a new trace $\sigma_{test}$ is collected from run-time monitoring, its diagnosis $d_{\sigma_{test}}$ is computed through online alignment-based conformance checking. Based on $d_{\sigma_{test}}$, the trace is assigned to the closest cluster and consequently labeled as a $comp_{i}$ anomaly if the closest cluster was previously labeled as containing diagnosis tuples affected by a fault in the i-th component. This completes the detection and explanation step, which uses diagnoses (non-conformances) of a trace with respect to its expected process flow to detect anomalies, diagnose their root cause, and localize their origin.
@@ 4.1 Behavior Characterization
In this step, we simulate the RBC/RBC Handover scenario through a high-level description of its process according to the SUBSET-026 document of the ERTMS standard. First, we compiled the various sequences identified in the document under the chosen scenario into a BPMN model. Since the model describes a workflow from start to end, the model can be simulated, and various traces can be collected.
After the BPMN traces are collected, they are further pre-processed to mimic a realistic monitoring process of the computer-based system. Specifically, we aim to mimic the procedural programming paradigm, which is commonly found in the real implementation of ERTMS/ETCS systems. Using the activities of the traces generated through the BPMN model, our approach records the invocation of a procedure each time a new activity is executed as follows.
Let $X$ be an ERTMS/ETCS component and $\mathcal{P}_{X}$ be the procedures implemented by $X$. One of the procedures is executed with a given probability for each distinct activity carried out by the component. It is expected that normal behavior follows a given control-flow with a high probability, whereas anomalous behavior involves the alteration of such normal control flow with the improper execution of other procedures. This is enforced using the following probability scheme. Let $\mathcal{A}_{HO}$ be the set of activities of the RBC/RBC Handover scenario, $\mathcal{X}=\{ARBC,HRBC,EVC,RTM\}$ the set of ERTMS/ETCS components involved in the use case, $\mathcal{P}_{a,X}=\{p_{0},\dots,p_{q_{X}}\}$ the $q_{X}\in\mathbb{N}$ procedures implemented by $X\in\mathcal{X}$, and $\rho\in[0,1)$ a real number. We define a procedure execution probability scheme $PEP$ for each $\mathcal{P}_{a,X}$ associated with $a\in\mathcal{A}_{HO}$ that assigns probability values to the procedures so that one procedure is much more likely to be chosen than others, allowing the simulation of traces that are very similar to each other. In particular, let us consider a real number $\rho\in\mathbb{R}$ and the following function definition:
$PEP:\mathbb{R}\times(\mathcal{P}_{a,X})^{q_{x}}\rightarrow\mathbb{R}^{q_{x}},$
$PEP(\rho,p_{0},\dots,p_{q_{X}})=(pr_{p_{0}},pr_{p_{1}},\dots,pr_{p_{q_{X}}})$, where $pr_{p_{0}}=\rho$ and the other probabilities are equal to $\frac{1-\rho}{q_{X}-1}$. For example, given $\mathcal{P}_{a,X}=\{p_{0},\dots,p_{q_{X}}\}$, if $\rho=0.99$, the application of $PEP$ to $\mathcal{P}_{a,X}$ is such that $p_{0}$ is selected with 99% probability, whereas there is a 0.1% probability that other procedures are executed.
Following the procedure above, we generate the execution traces of RBC/RBC Handover as follows.
@@
    Set $q_{X}$ for each component $X$ involved in the RBC/RBC Handover use case;
@@
    For each activity $a\in\mathcal{A}_{HO}$, enforce a random order to the procedures that may be executed for that activity, i.e., generate $\mathcal{P}_{a,X}$;
@@
    For each activity of the $N_{norm}$ traces, sample a procedure through $PEP$ using $\rho$.
In this preliminary setup, we have decided to set $q_{X}=10$ for each component and generate $N_{norm}=100$ traces. The dataset generation procedure has been published on the GitHub repository [^3] https://github.com/francescovitale/pm_ertms.
@@
We used the integer linear programming-based miner with 75% tolerance to noise to build a workflow Petri net from the $N_{norm}$ traces. Fig. 4 shows the Petri net that the algorithm extracted. The Petri net contains two specific places: the source and the sink. The source place, marked with ‘1’, allows triggering the first transition, whereas the sink place marks the end of the workflow. The red boxes outline those places with more than one outgoing arc, which means that in the presence of a single token in the place, only one of the outgoing paths can be followed — an exclusive control-flow pattern. The blue boxes outline those transitions with more than one outgoing arc, which means that multiple concurrent paths are activated when the transition is triggered — a concurrent control-flow pattern. The Petri net has 21 places, 20 transitions and 78 arcs. The Woflan diagnosis [verbeek2001woflan] reports that there always exists a path leading to the sink and that there are no dead transitions.
@@ 4.2 Behavior Diagnosis
In this part, we aim to extract online actions from run-time traces and check them against the execution model obtained in the previous step. We can simulate run-time traces with the same procedure above. In particular, we simulated $N_{anom}=400$ anomalous traces and injected control-flow anomalies. As there are 4 components, we split the $N_{anom}$ traces into four sets of 100 traces, which we refer to as $N_{anom,ARBC}$ traces, $N_{anom,HRBC}$ traces, $N_{anom,EVC}$ traces, and $N_{anom,RTM}$ traces. Each set of traces is injected with a different type of anomaly. The injection procedure is as follows. Let $X$ be one of the four components and let us consider its corresponding set of $N_{anom,X}$ traces. We scan the traces procedure-by-procedure and, if the procedure is executed by $X$, we randomly inject either a wrongly-ordered, skipped or wrong procedure control-flow anomaly. To account for the randomness of the simulation, we replicated the simulation ten times with different random seeds.
After we obtained the four sets of traces, we proceeded with the extraction of online actions by splitting each trace into multiple traces according to a window size. The number of online actions during this phase depends on the window size. We set three different window sizes, 5, 10 and 15, to evaluate the change in the results of the subsequent steps. For example, if a given trace has 20 traces and the window size is 10, it is split into 2 subtraces, where the first subtrace contains the first 10 online actions, whereas the second subtrace is the entire trace. Regardless of the window size, each subtrace is checked against the execution model to collect the conformance checking diagnoses to use in the subsequent steps.
@@ 4.3 Anomaly Detection and Explanation
The conformance checking diagnoses of the four sets of traces are clustered to identify the different anomalies. In particular, to evaluate the quality of different clustering techniques in correctly classifying the anomalies, we split the diagnoses into a training and a test set. We held out 25% of the diagnoses and used the training set to extract the clusters. Hence, there are 300 training diagnoses and 100 test diagnoses. Each cluster obtained from the training set was labeled with the type of anomaly injected. For example, all the clusters formed with the diagnoses of $N_{anom,ARBC}$ traces are labeled ARBC. The test diagnoses are subsequently assigned to each cluster and labeled accordingly. For completeness of our analysis, we selected multiple clustering algorithms, ensuring to select candidate clustering algorithms that are as diverse as possible between each other. We selected clustering algorithms that rely (e.g., K-Means) or do not rely on centroids (e.g., DBSCAN), algorithms that automatically derive the optimal number of clusters, and some that require the user to input the desired number of clusters as a parameter. Overall, we selected 10 algorithms: BIRCH, K-Means, Mini-Batch KMeans, Ward Hierarchical Clustering, Spectral Clustering, HDBSCAN, DBSCAN, OPTICS, MeanShift, AffinityPropagation, whose implementation is available in the open-source Python library scikit-learn. Algorithms that required the number of clusters as input were exercised 5 times, each with the input parameter of 10, 30, and 50 clusters.
The conformance checking diagnoses collect the misalignments of the 40 possible procedures. These procedures are categorized according to the component that executes them: procedures 0 to 9: ARBC component; procedures 10 to 19: EVC component; procedures 20 to 29: HRBC component; procedures 30 to 39: RTM component, thus $comp=\{ARBC,EVC,HRBC,RTM\}$. For each cluster, the anomaly explanation $s_{c}$ will be used to compute the quantities $s_{c,ARBC}$, $s_{c,EVC}$, $s_{c,HRBC}$, $s_{c,RTM}$, each quantifying the per-activity misalignment. Consequently, and regardless of the specific clustering algorithm, each cluster is labeled according to the process in Section 3.3.3 either as ARBC, EVC, HRBC or RTM.
@@ 5.1 Evaluation Metrics
There is a wide variety of metrics to measure the classification performance of multi-class classifiers [hossin2015review], the most common being accuracy, typically used in its balanced formulation. Other commonly used metrics are those that are primarily meant for binary classification (e.g., precision, recall, F-measure, matthews correlation coefficient) but may be adapted to multi-class scenarios by performing an average per class. For this study, we choose the balanced accuracy to quantify the performance of clustering algorithms.
In addition to the classification metrics, there is a subset of metrics that aim at assessing the performance of clustering processes. These metrics stem from the idea that a cluster should contain all and only instances of a specific class. To evaluate such capability, two metrics are usually employed: homogeneity, which determines whether a cluster contains coherent data points, and completeness, which evaluates whether data points of a given class are elements of the same cluster. The V-measure [rosenberg2007v] computes the harmonic mean of homogeneity and completeness. In this study, we use the V-measure to evaluate the quality of the clustering process.
Finally, we evaluate the effectiveness of the anomaly explanation mechanism discussed in Section 3.3.2. However, in this case, instead of computing the per-cluster explanation, we consider the explanation of each trace in the test set individually. Then, given a specific component, we sum all the test trace explanations, obtaining the four global anomaly explanations $S_{ARBC}$, $S_{EVC}$, $S_{HRBC}$ and $S_{RTM}$.
5.2 Anomaly Detection Results